High Country, Inc., produces and sells many recreational products. The company has just opened a new plant to produce a folding camp cot that will be marketed throughout the United States. The following cost and revenue data relate to May, the first month of the plant’s operation:

Beginning inventory 0
Units produced 35,000
Units sold 30,000
Selling price per unit $ 75
Selling and administrative expenses:
Variable per unit $ 3
Fixed (per month) $ 559,000
Manufacturing costs:
Direct materials cost per unit $ 15
Direct labor cost per unit $ 10
Variable manufacturing overhead cost per unit $ 4
Fixed manufacturing overhead cost (per month) $ 560,000

Management is anxious to assess the profitability of the new camp cot during the month of May.

Required:

1. Assume that the company uses absorption costing.

a. Determine the unit product cost.

b. Prepare an income statement for May.

2. Assume that the company uses variable costing.

a. Determine the unit product cost.

b. Prepare a contribution format income statement for May.
Answer #1
1a
Direct materials cost 15
Direct labor cost 10
Variable manufacturing overhead 4
Fixed manufacturing overhead 16 =560000/35000
Unit product cost 45
1b
Absorption Costing Income Statement
Sales 2250000 =30000*75
Cost of goods sold 1350000 =30000*45
Gross profit 900000
Selling and administrative expenses 649000 =559000+(30000*3)
Net operating income 251000
2a
Direct materials cost 15
Direct labor cost 10
Variable manufacturing overhead 4
Unit product cost 29
2b
Variable Costing Income Statement
Sales 2250000 =30000*75
Variable expenses:
Variable cost of goods sold 870000 =30000*29
Variable selling and administrative expenses 90000 =30000*3
960000
Contribution margin 1290000
Fixed expenses:
Fixed manufacturing overhead 560000
Fixed selling and administrative expenses 559000
1119000
Net operating income 171000